# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ Abstract.
Online advertising is a cornerstone of the Internet ecosystem, with advertising measurement playing a crucial role in optimizing efficiency. Ad measurement entails attributing desired behaviors, such as purchases, to ad exposures across various platforms, necessitating the collection of user activities across these platforms. As this practice faces increasing restrictions due to rising privacy concerns, safeguarding user privacy in this context is imperative. Our work is the first to formulate the real-world challenge of advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduce Ads-BPC, a novel user-level differential privacy protection scheme for advertising measurement results. This approach optimizes global noise power and results in a non-identically distributed noise distribution that preserves differential privacy while enhancing measurement accuracy. Through experiments on both real-world advertising campaigns and synthetic datasets, Ads-BPC achieves a 25% to 50% increase in accuracy over existing streaming DP mechanisms applied to advertising measurement. This highlights our method’s effectiveness in achieving superior accuracy alongside a formal privacy guarantee, thereby advancing the state-of-the-art in privacy-preserving advertising measurement.
Differential Privacy, User Level Privacy, Advertising Measurement, Impression, Conversion, Multi-touch Attribution, Online Streaming Dataset
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
@@
Online advertising is essential in the Internet ecosystem because it fuels the economy of the web and enables businesses to connect with their target audience effectively. Advertising measurement systems (yun2020challenges, ; guha2010challenges, ; briggs1997advertising, ; yuan2013real, ) are pivotal in the online advertising landscape. They aim to assess the effectiveness of advertising campaigns by determining the total number of conversions—such as purchases, sign-ups, downloads, or other desired actions—that occur as a direct result of ad exposure. To reach a more diverse audience, advertisers often set up ad campaigns across various platforms, including Google Ads, Facebook Ads, and other digital advertising networks. To effectively monitor the impact of these ads on each platform, appropriate tracking mechanisms are in place to allow for real-time monitoring of advertising campaigns across the selected channels.
@@
The ads measurement process encompasses two major steps: attribution and computation. The attribution step involves analyzing the customer’s online journey, starting from initial ad exposure to the ultimate conversion. Attribution aims to identify which exposures (touchpoints) contributed to conversions. The method that allocates credit to various advertising channels or touchpoints throughout the customer journey is referred to as multi-touch attribution (MTA). MTA recognizes the multiple interactions a customer may have with different advertising channels before making a conversion. By distributing credit among these touchpoints based on their influence on the conversion, MTA models offer a more nuanced understanding of each channel’s effectiveness within the advertising mix. The computation step is to calculate the total conversion value based on the attribution results to gain insights into the effectiveness of different advertising channels. Based on the such insights, advertisers can improve their advertising campaigns by reallocating budget, adjusting targeting parameters, refining ad designs, or tweaking other campaign elements to improve their return on investment and user advertising experience.
@@
Ad measurement requires collecting a variety of user data, including the ads users have viewed, clicks registered by the ad platform, and conversion data from the advertiser’s end. Traditionally, this involves automatically tracking user behavior across both ad publishers’ platforms and advertisers’ platforms, utilizing tools such as pixels, tags, or software development kits (SDKs). However, the tracking and collection of user data across sites have raised increasing privacy concerns regarding user consent, data security, and the potential misuse of personal information (toubiana2010adnostic, ; yun2020challenges, ; haddadi2011targeted, ; ullah2020privacy, ). These concerns have led to legislative actions, such as the General Data Protection Regulation (GDPR) in Europe (voigt2017eu, ) and industry initiatives such as Apple’s App Tracking Transparency (AppleATT, ), which impose varying degrees of restrictions on cross-site tracking. Recent studies have developed privacy-preserving mechanisms to support advertising measurement (ayala2022show, ; pfeiffer2021masked, ; rosales2012post, ). However, these approaches lack formal privacy guarantees. Our work addresses this gap by introducing a formal differential privacy guarantee for advertising measurement outcomes.
@@
Differential privacy (dwork06Calibrating, ; dworkKMM06:ourdata, ; zcdp, ; renyidp, ; fdp, ) provides a mathematical framework for adding carefully calibrated noise to query results or datasets to protect individual data record while still allowing for meaningful data analysis. The calibrated noise present makes it nearly impossible to discern whether the dataset contains a specific data record, thereby safeguarding sensitive information. Differential privacy has become a key component in modern privacy protection efforts in research areas such as outsourced database systems (near2021differential, ), machine learning model training data protection (abadi2016deep, ), healthcare (dankar2013practicing, ), census statistics (dwork2010differential, ; abowd2018uscensus, ), and data collection (erlingsson2014rappor, ; feldman2023stronger, ).
@@
In this paper, we undertake the novel task of designing a differentially private mechanism to protect individual user privacy in MTA ad measurement systems. Employing differential privacy in such practical settings presents two challenges. First, maintaining a continuous data stream poses significant hurdles for the effective implementation of differential privacy. In advertising campaigns, measurement results must be continuously disclosed to both advertisers and ad providers in real-time, per time slot (e.g., daily or hourly). This is crucial as it allows advertisers to observe the outcomes of their campaigns and adjust their strategies accordingly. Unlike one-time data releases in database queries, continuous streams gradually consume the differential privacy budget, leading to significant accuracy deterioration and making the analysis results less valuable from a business standpoint.
@@
Second, MTA involves records from various platforms, making traditional event-level differential privacy insufficient. In differential privacy, the granularity of privacy refers to the level at which DP guarantees are provided, indicating the unit of data protected by DP. Event-level DP protects whether a specific user action (such as clicking an ad or making a purchase) is included in the dataset, whereas user-level DP provides privacy protection at the level of individual users, covering all event-level records such as views, clicks, and conversions of a particular user. Consequently, user-level DP protects whether a user engaged with the ad campaign at any point, offering a stronger and more meaningful level of privacy protection than event-level DP. For instance, protecting whether a user has ever engaged with an ad for diabetic medication is more meaningful than merely protecting whether a user has clicked on an Instagram ad for diabetic medication.
In our pursuit of providing stronger DP protection for users in practical online advertising systems, this paper presents the following contributions:
  1. (1)
@@
    We are the first to introduce and formalize the problem regarding MTA advertising measurement systems, including both attribution and computation steps, which require the release of measurement results on a streaming basis.
  2. (2)
@@
    We devised a user-level streaming differentially private algorithm, Ads-BPC, tailored to address the challenges in MTA advertising measurement. Mathematically, we formulate the noise calibration process as a total noise variance minimization problem and derive a closed-form solution that adheres to a predetermined privacy budget. This non-identically distributed noise distribution preserves differential privacy and increases measurement accuracy.
  3. (3)
@@
    Through comprehensive experiments with both synthetic and real-world datasets, our algorithm achieves notable improvements in accuracy, ranging from 25% to 50%, compared to previous works. Our results indicate that our algorithm can provide more accurate and meaningful insights for advertisers while ensuring users’ privacy.
@@
This paper is structured as follows: Section 2 illustrates how advertising measurement operates through concrete examples and discuss related concepts in differential privacy. Section 3 explores previous work in the differential privacy area. Section 4 formally defines our user-level differentially private advertising measurement system, formulates it as a convex optimization problem, and presents an efficient algorithm to address it. Section 5 presents discussions on the proposed method. Finally, Section 6 presents the results of comprehensive experiments conducted on both synthetic and real-world datasets using our algorithm.
2. Ads Measurement and Differential Privacy
@@
In this section, we elucidate advertising measurement with multi-touch attribution as well as differential privacy definitions in the advertising measurement context. Mathematical notations used throughout this paper are summarized in Table 1.
Table 1. Table of Notation
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$: $k$-th impression of user $u$.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$: $k$-th conversion of user $u$.
@@
$\mathbf{W}$: Prefix-sum matrix.
@@
2.1. An Example of Advertising Measurement
@@
Impression Events. An impression event encompasses several key elements that define the context of an ad being served to a user. These elements include: 1) User ID, 2) Publisher ID, 3) Advertiser ID, 4) Timestamp, and 5) Metadata (supplementary information related to the impression, such as the type of user interaction (click or view) and ad format, providing the context for further analysis). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$ to express the $k$-th impression of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k})$ to express the metadata of the impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,k}$. In the case where an advertiser has multiple products, we use different Advertiser ID for different products. Table 2 shows an example of the impression table, where each row represents an impression event of users interacting (viewing or clicking) with ads on various publishers.
Table 2. Impression Table I on Day 1
@@
Conversion Events. Conversion events represent pivotal moments in the advertising pipeline when users generate value for advertisers after interacting with ads on the publishers’ platforms. These events encompass 1) User ID, 2) Advertiser ID, 3) Timestamp, and 4) Metadata: Additional data that characterize the conversion, including its type (e.g., purchase or sign-up) and value (e.g., purchase amount). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$ to express the $k$-th conversion of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k})$ to express the metadata of the conversion ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,k}$. Table 3 shows an example of the conversion table, where each row represents a conversion, such as a purchase of Nike shoes.
Table 3. Conversion Table C on Day 1
@@
Advertising Measurement Process The gist of ad measurement can be distilled into a two-step process. In the first step, conversions are attributed to one or more impressions using predefined logic or attribution models. This step is referred to as “attribution,” determining which impressions played a role in driving the eventual conversion event. Section 2.1.1 gives more details about attribution logic. Subsequently, in the second step, queries are executed on the resulting attributed dataset to extract actionable insights. These queries may involve counting the total number of conversions attributed to a subset of impressions, calculating the corresponding conversion rate, or assessing the total return on ad spend for that particular subset.
Table 4. Table after Joining Impression Table I and Conversion Table C for Day 1 based on User-ID and Ad-ID
@@
2.1.1. Attribution Logic
For attribution, we assume there is a trusted third party that joins the impression table and the conversion table. Listing 1 shows a simplified SQL query that joins these two tables. Table 4 shows the result of joining Tables 2 and 3. Then the third party assigns the attribution weight of a conversion to each relevant impression. The following are several commonly used attribution logic.

### v2

@@ Abstract
Online advertising is a cornerstone of the Internet ecosystem, with advertising measurement playing a crucial role in optimizing efficiency. Ad measurement entails attributing desired behaviors, such as purchases, to ad exposures across various platforms, necessitating the collection of user activities across these platforms. As this practice faces increasing restrictions due to rising privacy concerns, safeguarding user privacy in this context is imperative. Our work is the first to formulate the real-world challenge of advertising measurement systems with real-time reporting of streaming data in advertising campaigns. We introduce AdsBPC, a novel user-level differential privacy protection scheme for online advertising measurement results. This approach optimizes global noise power and results in a non-identically distributed noise distribution that preserves differential privacy while enhancing measurement accuracy. Through experiments on both real-world advertising campaigns and synthetic datasets, AdsBPC achieves a 33% to 95% increase in accuracy over existing streaming DP mechanisms applied to advertising measurement. This highlights our method’s effectiveness in achieving superior accuracy alongside a formal privacy guarantee, thereby advancing the state-of-the-art in privacy-preserving advertising measurement.
1 Introduction
@@
Online advertising is essential in the Internet ecosystem as it fuels the economy of the web and enables businesses to connect with their target audience effectively. Advertising measurement systems [55, 23, 7, 54] are pivotal in the online advertising landscape. They aim to assess the effectiveness of advertising campaigns by determining the total number of conversions—such as purchases, sign-ups, downloads, or other desired actions—that occur as a result of ad exposure. To reach a more diverse audience, advertisers often set up ad campaigns across various platforms, including Google Ads, Facebook Ads, and other digital advertising networks. To effectively gauge the impact of these ads on each platform, tracking mechanisms are in place to monitor advertising campaigns across the selected channels.
@@
The ad measurement process encompasses two major steps: attribution and computation. The attribution step involves analyzing the customer’s online journey, starting from initial ad exposure to the ultimate conversion. Attribution aims to identify which exposures (touchpoints) contributed to conversions. The method that allocates credit to various advertising channels or touchpoints throughout the customer journey is referred to as multi-touch attribution (MTA). MTA recognizes the multiple interactions a customer may have with different advertising channels before making a conversion. By distributing credit among these touchpoints based on their influence on the conversion, MTA models offer a more nuanced understanding of each channel’s effectiveness within the advertising mix. The computation step is to calculate the total conversion value based on the attribution results to gain insights into the effectiveness of different advertising channels. Based on the such insights, advertisers can improve their advertising campaigns by reallocating budget, adjusting targeting parameters, refining ad designs, or tweaking other campaign elements to improve their return on investment and users’ advertising experience.
@@
Ad measurement requires collecting a variety of user data, including the ads users have viewed, clicks registered by the ad platform, and conversion data from the advertiser’s end. Traditionally, this involves automatically tracking user behavior across both ad publishers’ platforms and advertisers’ platforms, utilizing tools such as pixels, tags, or software development kits (SDKs). However,the tracking and collection of user data across sites have raised increasing privacy concerns regarding user consent, data security, and the potential misuse of personal information [46, 55, 25, 47]. These privacy concerns have led to legislative actions, such as the General Data Protection Regulation (GDPR) in Europe [48] and industry initiatives such as Apple’s App Tracking Transparency [5], which impose varying degrees of restrictions on cross-site tracking. Recent studies have developed privacy-preserving mechanisms to support advertising measurement [6, 37, 39]. However, these approaches lack formal privacy guarantees. Our work addresses this gap by introducing a formal differential privacy guarantee for advertising measurement outcomes.
@@
Differential privacy [20, 8] provides a mathematical framework for adding carefully calibrated noise to query results or datasets to protect individual data record while still allowing for meaningful data analysis. The calibrated noise makes it nearly impossible to discern whether the dataset contains a specific data record or not, thereby safeguarding the privacy of individual data record. Differential privacy has become a key component in modern privacy protection efforts in research areas such as database systems [36], machine learning model training data protection [3], and census statistics [4].
@@
Employing a differentially private mechanism to protect individual user privacy in a practical MTA ad measurement setting presents two challenges. First, maintaining a continuous data stream poses significant hurdles for the effective implementation of differential privacy. In advertising campaigns, measurement results must be continuously disclosed to both advertisers and ad providers in real-time, per time slot (e.g., daily or hourly). This is crucial as it allows advertisers to observe the outcomes of their campaigns and adjust their strategies accordingly. Unlike one-time data releases in database queries, continuous streams gradually consume the differential privacy budget, leading to significant accuracy deterioration and making the analysis results less valuable from a business standpoint. Existing differentially private streaming algorithms [11, 21, 19] primarily focus on protecting prefix-sum queries (i.e., releasing the cumulative sum up to the current day) and minimizing the root mean squared error. These algorithms are limited in the types of supported queries and objective functions.
@@
Second, MTA involves records from various platforms, making traditional event-level differential privacy insufficient. In differential privacy, the granularity of privacy refers to the level at which DP guarantees are provided, indicating the unit of data protected by DP. Event-level DP protects whether a specific user action (such as clicking an ad or making a purchase) is included in the dataset, whereas user-level DP provides privacy protection at the level of individual users, covering all event-level records such as views, clicks, and conversions of a particular user. Consequently, user-level DP protects whether a user engaged with the ad campaign at any point, offering a stronger and more meaningful level of privacy protection than event-level DP. For instance, protecting whether a user has ever engaged with an ad campaign for diabetic medication is more meaningful than merely protecting whether a user has clicked on an Instagram ad for diabetic medication.
To address the two challenges above, we propose the AdsBPC algorithm, which offers greater flexibility by supporting various types of online queries and objective functions. This is achieved without significant modifications to the algorithm’s structure—only the initialization of the noise scale parameters needs to be adjusted. The algorithm can then adaptively produce the desired results based on the provided data. Our AdsBPC algorithm is also designed to provide a formal user-level differential privacy. The primary contributions of this paper are as follows:
  1. 1.
@@
    We are the first to introduce and formalize the problem regarding online MTA advertising measurement systems, including both attribution and computation steps, which require the release of measurement results on a streaming basis.
  2. 2.
@@
    We devised a user-level streaming differentially private algorithm, AdsBPC, tailored to address the challenges in MTA advertising measurement. Mathematically, we formulate the noise calibration process as a noise variance minimization problem with privacy budget constraints.
  3. 3.
    We developed a differentially private algorithm designed to accurately bound per-day user contributions. This approach reduces the sensitivity of AdsBPC, thereby enhancing accuracy while maintaining a fixed privacy budget.
  4. 4.
@@
    Through comprehensive experiments with both synthetic and real-world datasets, our algorithm achieves notable improvements in accuracy, ranging from 33% to 95%, compared to previous works. Our results indicate that our algorithm can provide more accurate and meaningful insights for advertisers while ensuring users’ privacy.
@@
This paper is structured as follows: Section 2 illustrates how advertising measurement operates through concrete examples and discuss related concepts in differential privacy. Section 3 explores previous work in the differential privacy area. Section 4 formally defines our user-level differentially private advertising measurement system and presents an efficient algorithm to address it. Section 5 presents discussions on the proposed method. Finally, Section 6 presents the results of comprehensive experiments conducted on both synthetic and real-world datasets using our algorithm.
2 Ad Measurement and Differential Privacy
@@
In this section, we elucidate advertising measurement with multi-touch attribution as well as differential privacy definitions in the advertising measurement context. Mathematical notations used throughout this paper are summarized in Table I.
TABLE I: Table of Notation
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,j}$: $j$-th impression of user $u$.
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,j}$: $j$-th conversion of user $u$.
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}$: The $j$-th query vector.
$\mathbf{Q}$: Query workload matrix, it’s a stack of query vectors.
@@
$r_{i}$: User contribution bound on day $i$.
$\sigma_{i}$: Noise scale on day $i$.
$l$: Number of days using differentially private quantile.
$p$: The quantile percentage used in the private quantile.
$s^{\uparrow},s^{\downarrow}$: Scaling parameters for updating the user contribution bound.
$T^{\uparrow},T^{\downarrow}$: Thresholds used in the sparse vector technique.
2.1 An Example of Advertising Measurement
@@
Impression Events. An impression event encompasses several key elements that define the context of an ad being served to a user. These elements include: 1) User ID, 2) Publisher ID, 3) Advertiser ID, 4) Timestamp, and 5) Metadata (supplementary information related to the impression, such as the type of user interaction (click or view) and ad format, providing the context for further analysis). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,j}$ to express the $j$-th impression of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,j})$ to express the metadata of the impression ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u,j}$. In the case where an advertiser has multiple products, we use different Advertiser IDs for different products. Table II shows an example of the impression table, where each row represents an impression event of users interacting (viewing or clicking) with ads on various publishers.
TABLE II: Impression Table I on Day 1
@@
Conversion Events Conversion events represent pivotal moments in the advertising pipeline when users generate value for advertisers after interacting with ads on the publishers’ platforms. These events encompass 1) User ID, 2) Advertiser ID, 3) Timestamp, and 4) Metadata: Additional data that characterize the conversion, including its type (e.g., purchase or sign-up) and value (e.g., purchase amount). We use the notation ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,j}$ to express the $j$-th conversion of user $u$, and use ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{\phi}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,j})$ to express the metadata of the conversion ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u,j}$. Table III shows an example of the conversion table, where each row represents a conversion, such as a purchase of Nike shoes.
TABLE III: Conversion Table C on Day 1
@@
Advertising Measurement Process The gist of ad measurement can be distilled into a two-step process. In the first step, conversions are attributed to one or more impressions using predefined logic or attribution models. This step is referred to as “attribution,” determining which impressions played a role in driving the eventual conversion event. Section 2.1.1 gives more details about attribution logic. Subsequently, in the second step, queries are executed on the resulting attributed dataset to derive valuable aggregate statistics. These queries may involve counting the total number of conversions attributed to a subset of impressions, calculating the corresponding conversion rate, or assessing the total return on an ad.
TABLE IV: Table after Joining Impression Table I and Conversion Table C for Day 1 based on User-ID and Ad-ID
@@
2.1.1 Attribution Logic
For attribution, we assume there is a trusted third party that joins the impression table and the conversion table. Listing 1 shows a simplified SQL query that joins these two tables. Table IV shows the result of joining Tables 2 and 3. Then the third party assigns the attribution weight of a conversion to each relevant impression. The following are several commonly used attribution logic. Last-Touch Attribution (LTA), also known as Last-Click Attribution, assigns full credit for a conversion to the last interaction or touchpoint that directly led to the conversion. First-Touch Attribution (FTA), as the name suggests, attributes full credit for a conversion to the first interaction or touchpoint that introduced the customer to the brand or campaign. Uniform Attribution (UNI) is an attribution model in advertising measurement that evenly distributes credit for conversions across all touchpoints in a customer’s journey, providing equal weight to each interaction.
Table V shows an example of the above attribution logics, given the joined Table IV. Since only ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ is relevant to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},1}$, both LTA, FTA, and UNI will assign a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ (similarly for ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{1},1}$). In the case of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{u_{2},2}$, where both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ are relevant, LTA attributes a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$ as it occurs last. FTA assigns a weight of 1 to ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ since it happens first. UNI, on the other hand, evenly distributes the weight, assigning 0.5 to both ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},1}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{i}}}_{u_{2},2}$.
There are other more sophisticated attribution algorithms. For example, Incremental Value Heuristics (IVH) [41] measures the impact of an impression by comparing results between a group exposed to the ads and a control group that was not exposed. Shapley Value methods (SV) [57, 41] originate from game theory and allocate credit to each impression based on its marginal contribution. Data-Driven Attribution (DDA) [40, 26] uses machine learning models to analyze historical data and assign attribution weights by comparing counterfactual impression patterns to reveal how different touchpoints work together to drive conversions. Our Ads-BPC algorithm is compatible with these advanced attribution models, provided that the normalized total weight assigned to each publisher for any given conversion ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}$ does not exceed 1.